40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  PSL  Not logged in ELOG logo
Message ID: 2343     Entry time: Mon May 13 19:12:56 2019
Author: anchal 
Type: DailyProgress 
Category: Other 
Subject: Set experiment for 2.6 mW transmission from cavities. 

I tuned the intensity control and changed autolock parameters to get 2.6 mW light at the output of cavities on both paths. At present, the South Path has 74 % mode matching and North Path has 63% mode matching. These numbers were higher yesterday, so I'm suspecting something shifted over a day.

I also maximized gains on PMC loops and FSS loops:

  North South
PMC Gain (dB) 30 (Max) 21
FSS Common Gain (V) [32dB/V] 3.2 10 (Max)

In FSS loop, I kept gain margin of 0.64 dB (i.e. the oscillations are seen at 0.2V above the set value of Common Gain) and in PMC loop, I kept it at 2 dB. I yet have to characterize noise in the FSS error signal and see what I can make better. On another note, I did a rough calculation today and I think shotnoise intercept current for our RFPDs is ~2.6 mA which means atleast 3.45 mW light should be falling on them for them to be shotnoise limited. I have to check the validity of my rough estimation but if this is true, then our FSS RFPDs are not shot noise limited.

ELOG V3.1.3-